OpenStudy (anonymous):

Find the leg of each isosceles right triangle when the hypotenuse is of the given measure. Given 6(square root symbol)2

OpenStudy (anonymous):

Okay, what's the measure?

OpenStudy (anonymous):

a^2 + b^2 = c^2 a^2 = b^2 in an isosceles a^2 + a^2 2a^2 = c^2 for an isosceles.

OpenStudy (anonymous):

6(square root symbol)2